Find all complex eigenvalues of the matrix A= 1 + Drag and drop your files or click to browse...
Answer #1

A=\begin{pmatrix}-i&-4\\ 1&2i\end{pmatrix}

(A-xI) \:\:is,

\begin{pmatrix}-i&-4\\ 1&2i\end{pmatrix}-x\begin{pmatrix}1&0\\ 0&1\end{pmatrix}

=\begin{pmatrix}-i-x&-4\\ 1&2i-x\end{pmatrix}

take determinant =0

\left(-i-x\right)\left(2i-x\right)-\left(-4\right)\cdot \:1=0

-2i^2-2ix+ix+x^2+4=0

-2(-1)-ix+x^2+4=0

2-ix+x^2+4=0

x^2+2-ix+4=0

x^2-ix+6=0

x_{1,\:2}=\frac{-b\pm \sqrt{b^2-4ac}}{2a}

x_{1,\:2}=\frac{-\left(-i\right)\pm \:\sqrt{\left(-i\right)^2-4\left(1\right)\left(6\right)}}{2\left(1\right)}

x_{1,\:2}=\frac{i\pm \:\sqrt{-1-24}}{2}

x_{1,\:2}=\frac{i\pm \:\sqrt{-25}}{2}

x_{1,\:2}=\frac{i\pm \:5i}{2}

x=\frac{i+5i}{2},\:\:\:\:\:\:\:\:\:\:\:\:x=\frac{i-5i}{2}

x=\frac{6i}{2},\:\:\:\:\:\:\:\:\:\:\:\:x=\frac{-4i}{2}

{\color{Red} x=3i,\:\:\:\:\:\:\:\:\:\:\:\:x=-2i}........................eigenvalues